Find all numbers for which the rational expression is not defined. Separate by comma if needed.
x^2-4/x^2+4x-21

Answer by stanbon(75887)

Find all numbers for which the rational expression is not defined. Separate by comma if needed.
x^2-4/x^2+4x-21
---
Not defined if x^2+4x-21 = 0
Factor:
x^2+7x-3x-21 = 0
x(x+7)-3(x+7) = 0
(x+7)(x-3) = 0
---
Undefined when x = -7 or x = 3
=================================
Cheers,
Stan H.
=======